# Flat-band FFLO State from Quantum Geometric Discrepancy

### v2

@@ Abstract
with $p_{\bm{k}}=\cos k_{x}+\cos k_{y}$. For brevity we set the lattice constant $a=1$ and $\mu_{\uparrow}=\mu_{\downarrow}=\mu$. With the spin index $\sigma=\uparrow,\downarrow$, the fermion operators are defined as $\hat{c}_{\bm{k}\sigma}=(\hat{c}_{A,\bm{k}\sigma},\hat{c}_{B,\bm{k}\sigma})^{T}$. The Pauli matrices $\lambda_{i}$ act on orbital space, and time-reversal flavors are related by $m_{\sigma=\uparrow/\downarrow}=\pm 1$. The periodic function $\zeta_{\sigma}p_{\bm{k}}$ contains spin-dependent long range hoppings [see Fig. 2(a)]. As plotted in Fig. 2(b), this model Hamiltonian possesses a pair of ideal isolated flat bands with band gap $\Delta_{\mathrm{gap}}=2t$. For each spin flavor, the band dispersion $\xi_{\bm{k}\sigma}$ as well as Bloch function $u_{\bm{k}\sigma}$ can be solved as $\xi_{\bm{k}\sigma}=\pm t-\mu$ and
@@
Given the Bloch function $u_{\bm{k}\sigma}$, the spin-dependent quantum metrics can be expressed as $\mathcal{G}_{ab}(\bm{k},\sigma)=\zeta_{\sigma}^{2}\sin(k_{a})\sin(k_{b})/4$. When considering electrons with opposite momentum $\bm{k}$ that yield spin-singlet pairing, the time-reversal partners $(\bm{k},\uparrow)$ and $(-\bm{k},\downarrow)$ form a Cooper pair. Crucially, $(\bm{k},\uparrow)$ and $(-\bm{k},\downarrow)$ are degenerate in energy as $\xi_{\bm{k}\uparrow}=\xi_{-\bm{k}\downarrow}$, but are polarized in quantum metrics when $\zeta_{\uparrow}\neq\zeta_{\downarrow}$, meaning $\mathcal{G}(\bm{k},\uparrow)\neq\mathcal{G}(-\bm{k},\downarrow)$. In this case, $\mathcal{T}$ symmetry of the system is explicitly broken, which is evidenced by the $u_{\bm{k}\uparrow}\neq u_{-\bm{k}\downarrow}^{*}$, rather than by an energy difference. For the $\zeta$-lattice model, we define the dimensionless parameter $\eta=|\zeta_{\uparrow}-\zeta_{\downarrow}|/(\zeta_{\uparrow}+\zeta_{\downarrow})$ as a global measure of QGD. In the following, we will illustrate the role of QGD in the formation of flat-band FFLO states.
@@
Figure 2: (a) The $\zeta$-lattice with spin-dependent long-range hoppings. $\zeta_{\sigma}$ denotes the quantum metric of spin-$\sigma$ sector. (b) The energy dispersions in flat-band limit ($W=0$) and finite bandwidth ($W\neq 0$). (c) $T_{c\bm{Q}}(\eta)$ (dashed blue), $T_{c0}(\eta)$ (solid red), and the corresponding $Q(\eta)$ (solid green) from Eq. (6) as a funtion of $\eta$. The orange dot denotes the Lifshitz point. The $\bm{Q}(\eta)$ calculated from the self-consistent mean-field (MF) theory is plotted in the dashed green line as a comparison. (d) Calculated $T_{c\bm{q}}(\eta)/T_{0}$ from Eq. (6) on $\bm{q}=q(\hat{x}+\hat{y}),\,q\in\left[0,\pi\right]$, for $\eta=0$ and $\eta=0.25$, respectively. In our calculations, we set $\zeta_{\downarrow}=3$ and $\zeta_{\uparrow}=\zeta_{\downarrow}(1+\eta)/(1-\eta)$.
QGD-stabilized flat-band FFLO state.—In terms of Eq. (6), in Fig. 2(c) we plot $T_{c\bm{Q}}(\eta)$ (dashed blue) as well as the corresponding $\bm{Q}(\eta)$ (solid green). As a comparison, $T_{c0}(\eta)$ (solid red) is also depicted. In the BCS phase where $\eta<\eta_{c}=0.22$, $T_{c\bm{Q}}(\eta)$ is identical to $T_{c0}(\eta)$, and $\bm{Q}(\eta)=0$. However, when $\eta>\eta_{c}$, $T_{c\bm{Q}}(\eta)\neq T_{c0}(\eta)$, and a sharp transition occurs in $\bm{Q}(\eta)$ [from $Q(\eta_{c}-)=0$ to $Q(\eta_{c}+)\approx 0.31\pi$], implying a transition from BCS to FFLO phase. The tri-critical point (orange dot) of BCS, FFLO, and the normal phase is the so-called Lifshitz point, which is located at $(\eta_{c}=0.22,T=0.57T_{0})$. We also find that $\bm{Q}$ is four-fold degenerate, which are located at ±$\pm Q(\hat{x}+\hat{y})$ and ±$\pm Q(-\hat{x}+\hat{y})$ sup , so in our calculations we take $\bm{Q}=Q(\hat{x}+\hat{y})$ throughout this work.
@@
To further understand the QGD-driven FFLO transition and the role of AQD, we plot the $T_{c\bm{q}}(\eta)$ evaluated from Eq. (6) in Fig. 2(d), for $\eta=0$ (pink curve) and $\eta=0.25$ (orange curve), respectively. The highest critical temperature is determined by the minimum of $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$, which corresponds to the maximum of $T_{c\bm{q}}$. As we mentioned before, $\mathcal{T}$ symmetry ($\eta=0$) ensures the favorable BCS pairing state ($Q=0$) because of the global maximum of $T_{c\bm{q}}$ always occurs at $q=0$ (dashed green circle on the pink curve). Note that a local maximum of $T_{c\bm{q}}(0)$ can be seen at nonzero $q$ in Fig. 2(d). As QGD ($\eta\neq 0$) decreases $T_{c0}(\eta)$ significantly (the green dashed circle on the orange curve), it is possible that a nonzero $Q$ can be stabilized. Upon increasing $\eta$, the local maximum finally transitions to global maxima (the cyan dashed circle on the orange curve) at $Q(\eta)\neq 0$, where a FFLO phase can naturally manifest. Here our calculations are based on the square lattice model, and in Supplemental Material Note III sup we also show a triangular lattice model, where our theory still applies.
@@
Self-consistent mean field calculations—The pair susceptibility analysis enables us to find FFLO instabilities in the single-band description. Here we numerically show, based on self-consistent mean field study Kitamura et al. (2022), there is a first-order phase transition from BCS to FFLO phase driven by QGD within the previously introduced $\zeta$-lattice model [Eq. (S23)]. Near the critical temperature, the numerical results perfectly match Eq. (6). The mean-field decoupling of the attractive Hubbard interaction yields orbital-dependent order parameters given by $\Delta_{i\alpha}=-U\langle\hat{c}_{i\alpha\downarrow}\hat{c}_{i\alpha\uparrow}\rangle=\Delta_{\alpha}(\bm{q})e^{i\bm{q}\cdot\bm{r}_{i}}$ with $\bm{q}$ index. In the canonical ensemble, the Helmholtz free energy density $F$ is written as
@@
with the order parameters $\hat{\Delta}_{\bm{q}}=\mathrm{diag}[\Delta_{A}(\bm{q}),\Delta_{B}(\bm{q})]$. $\hat{\Delta}_{\bm{q}}$ and $\mu_{\bm{q}}$ can be determined self-consistently by solving the gap and number equations, respectively. Substituting $\hat{\Delta}_{\bm{q}}$ and $\mu_{\bm{q}}$ into Eq. (9), we can obtain $F(\bm{q})$ for each $\bm{q}$ (see more details in Supplemental Material Note II sup ).
@@
The results from mean-field calculations are summarized in Fig. 3. As shown in Fig. 3(a), the BCS as well as FFLO phases are colored in light blue and yellow, respectively. $T_{c0}(\eta)$ and $T_{c\bm{Q}}(\eta)$ are plotted in the dashed black and the solid red curves. The most favored pairing momentum $\bm{Q}(\eta)$ is independent of the temperature, resulting in a straight line for the boundary between the BCS and FFLO phases, which implies that the phase transition can only be driven by QGD, rather than the thermal fluctuations. The calculated $T_{c0}(\eta)$, $T_{c\bm{Q}}(\eta)$ and $\bm{Q}(\eta)$ [$\bm{Q}(\eta)$ is also plotted in Fig. 2(a) to compare with the one evaluated from the band projection method] exhibits quantitatively good agreements with the corresponding analytical results in Fig. 2(c). To provide a clear picture of the phase transition, in Fig. 3(c) to (e), we show the landscapes of free energy $F(\bm{q})$ for the three points (blue, green, and pink dots) in Fig. 3(a). Initially, in the BCS region ($\eta<\eta_{c}$) [Fig. 3(c)], the global minimum of the free energy $F(\bm{q})$ is at $\bm{q}=0$, which is stable under weak fluctuations. Some local minima exist at $\bm{q}\neq 0$ but are metastable. When $\eta>\eta_{c}$ [Fig. 3(d)], $\bm{q}=0$ turns into a local minimum but remains a metastable state, while some $\bm{q}=\bm{Q}$ points become global minima (the model respects $C_{4}$ rotational symmetry, so the minima have four-fold degeneracy in the whole Brillouin zone, see Supplemental Material Note III sup ), the FFLO phase can manifest. This indicates that the QGD energetically penalizes BCS pairing more than FFLO pairing. A first-order phase transition from BCS to FFLO occurs near $\eta_{c}$, which is consistent with the results obtained from the band-projection method. As $\eta$ continues to increase [Fig. 3(e)], the local minimum at $\bm{q}=0$ eventually nearly disappears.
@@
Robustness at finite bandwidth.—The QGD-induced BCS-FFLO phase transition has been examined in the ideal flat-band case, both analytically and numerically. However, a natural concern arises about whether the theory will break down at finite band dispersion. To show the robustness of QGD-induced BCS-FFLO transition with finite band dispersion, we modify the Bloch Hamiltonian in Eq. (S23) as $\mathcal{H}^{\prime}_{\bm{k}\sigma}=\mathcal{H}_{\bm{k}\sigma}-W\left(\cos k_{x}+\cos k_{y}\right)\lambda_{0}/4$, where an additional nearest hopping term brings a finite bandwidth $W$ as shown in Fig. 2(b). In this case, the Bloch wave functions remain unchanged. We keep $W\ll U$ to ensure a narrow bandwidth compared to the interaction. In Fig. 3(b), the calculated phase diagram at $W/U=0.12$ ($U=0.4t$) is shown. Obviously, compared to the exactly flat band case, the transition boundary line is temperature-dependent and becomes curved due to the nonzero $W$, requiring a larger QGD to reach the FFLO region compared with Fig. 3(a), especially at lower temperatures. This fact indicates that increasing the temperature may also drive the BCS-FFLO phase transition (e.g., along the dashed pink line at $\eta=0.25$ in Fig. 3(b)). The reason is, in Eq. (5), the form factor gets Fermi-surface averaged for a dispersive band, which weakens the effects of QGD. And the thermal excitation energy $k_{B}T$ influences the Fermi-surface average in Eq. (5), reshaping the transition boundary line. On the other hand, for conventional superconductors with a hkighly dispersive band, we have $W\gg U\gg k_{B}T$ in the weak coupling regime, where the QGD-driven mechanism for the FFLO state may not be effective (more discussions, see Supplemental Material Note II sup ), which is consistent with our intuition.

### v3

@@ IV Model Hamiltonian
with $p_{\bm{k}}=\cos k_{x}+\cos k_{y}$. For brevity we set the lattice constant $a=1$ and $\mu_{\sigma}=\mu$. With the spin index $\sigma=\uparrow,\downarrow$, the fermion operators are defined as $\hat{c}_{\bm{k}\sigma}=(\hat{c}_{A,\bm{k}\sigma},\hat{c}_{B,\bm{k}\sigma})^{T}$. The Pauli matrices $\lambda_{i}$ act on orbital space, and time-reversal flavors are related by $m_{\sigma=\uparrow/\downarrow}=\pm 1$. The periodic function $\zeta_{\sigma}p_{\bm{k}}$ contains spin-dependent long range hoppings [see Fig. 2(a)]. As plotted in Fig. 2(b), this model Hamiltonian possesses a pair of ideal isolated flat bands with band gap $\Delta_{\mathrm{gap}}=2t$. For each spin flavor, the band dispersion $\xi_{\bm{k}\sigma}$ as well as Bloch function $u_{\bm{k}\sigma}$ can be solved as $\xi_{\bm{k}\sigma}=\pm t-\mu$ and
@@
Given the Bloch function $u_{\bm{k}\sigma}$, the spin-dependent quantum metrics can be obtained as $\mathcal{G}_{ab}(\bm{k},\sigma)=\zeta_{\sigma}^{2}\sin(k_{a})\sin(k_{b})/4$.
@@
When considering electrons with opposite momentum $\bm{k}$ that yield spin-singlet pairing, the time-reversal partners $(\bm{k},\uparrow)$ and $(-\bm{k},\downarrow)$ form a Cooper pair. Crucially, $(\bm{k},\uparrow)$ and $(-\bm{k},\downarrow)$ are degenerate in energy as $\xi_{\bm{k}\uparrow}=\xi_{-\bm{k}\downarrow}$, but are polarized in quantum metrics when $\zeta_{\uparrow}\neq\zeta_{\downarrow}$, meaning $\mathcal{G}(\bm{k},\uparrow)\neq\mathcal{G}(-\bm{k},\downarrow)$. In this case, $\mathcal{T}$ symmetry of the system is broken, which is evidenced by $u_{\bm{k}\uparrow}\neq u_{-\bm{k}\downarrow}^{*}$, rather than by energy difference. For the $\zeta$-lattice model, we define the dimensionless parameter $\eta=|\zeta_{\uparrow}-\zeta_{\downarrow}|/(\zeta_{\uparrow}+\zeta_{\downarrow})$ as a global measure of QGD. In the following, we will illustrate the role of QGD in the formation of flat-band FFLO states.
V QGD-stabilized flat-band FFLO state
[…]
@@ V QGD-stabilized flat-band FFLO state
Figure 3: (a) $T_{c\bm{Q}}(\eta)$ (dashed blue), $T_{c0}(\eta)$ (solid red), and the corresponding $Q(\eta)$ (solid green) calculated from Eq. (6) as a funtion of $\eta$. The orange dot labels the Lifshitz point. For comparison, the dashed green line shows $Q(\eta)$ calculated from the self-consistent mean-field (MF) theory. (b) $T_{c\bm{q}}(\eta)/T_{0}$ from Eq. (6) for $\bm{q}=q(\hat{x}+\hat{y}),\,q\in\left[0,\pi\right]$, at $\eta=0$ and $\eta=0.25$, respectively. In our calculations, we set $\zeta_{\downarrow}=3$ and $\zeta_{\uparrow}=\zeta_{\downarrow}(1+\eta)/(1-\eta)$.
@@
To further understand the QGD-driven FFLO transition and the role of AQD, we plot the $T_{c\bm{q}}(\eta)$ evaluated from Eq. (6) in Fig. 3(b), for $\eta=0$ (pink curve) and $\eta=0.25$ (orange curve), respectively. The highest critical temperature is determined by the minimum of $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$, which corresponds to the maximum of $T_{c\bm{q}}$. As we mentioned before, $\mathcal{T}$ symmetry ($\eta=0$) ensures the favorable BCS pairing state ($Q=0$) because the global maximum of $T_{c\bm{q}}$ always occurs at $q=0$ (dashed green circle on the pink curve). Note that a local maximum of $T_{c\bm{q}}(0)$ can be seen at nonzero $q$ in Fig. 3(b). As QGD ($\eta\neq 0$) decreases $T_{c0}(\eta)$ significantly (the green dashed circle on the orange curve), it is possible that a nonzero $Q$ can be stabilized. Upon increasing $\eta$, the local maximum finally transitions to global maxima (the cyan dashed circle on the orange curve) at $Q(\eta)\neq 0$, where a FFLO phase can naturally manifest. Here our calculations are based on the square lattice model, and in SI III.A sup we also examine a triangular lattice model, where our theory still applies. Our analysis based on pairing susceptibility is supported by a recent follow-up work Zhang et al. (2025) through numerical determinantal quantum Monte Carlo calculations.
VI Self-consistent mean field calculations
@@
We have seen that the pairing susceptibility analysis, including Eq. (5) and (6), have a profound impact on identifying FFLO instabilities within the single-band description. To systematically verify these findings and get a full phase diagram, we perform numerical simulations using self-consistent mean-field theory Kitamura et al. (2022). These calculations clearly demonstrate a first-order BCS to FFLO phase transition, directly driven by QGD in our model Hamiltonian [Eq. (S30)]. For clarity, the mean-field decoupling of the attractive Hubbard interaction yields orbital-dependent order parameters as $\Delta_{i\alpha}=-U\langle\hat{c}_{i\alpha\downarrow}\hat{c}_{i\alpha\uparrow}\rangle=\Delta_{\alpha}(\bm{q})e^{i\bm{q}\cdot\bm{r}_{i}}$ with $\bm{q}$ index. In the canonical ensemble, the Helmholtz free energy density $F$ can be written as
@@
with the order parameters $\hat{\Delta}_{\bm{q}}=\mathrm{diag}[\Delta_{A}(\bm{q}),\Delta_{B}(\bm{q})]$. $\hat{\Delta}_{\bm{q}}$ and $\mu_{\bm{q}}$ can be determined self-consistently by solving the gap and number equations, respectively. Substituting $\hat{\Delta}_{\bm{q}}$ and $\mu_{\bm{q}}$ into Eq. (9), we can obtain $F(\bm{q})$ for each $\bm{q}$ (see more details in SI II sup ).
@@
The numerical results are summarized in Fig. 4. As shown in Fig. 4(a), the BCS and FFLO phases are colored light blue and yellow, respectively. The critical temperatures $T_{c0}(\eta)$ as well as $T_{c\bm{Q}}(\eta)$ are plotted in the dashed black and the solid red curves. We find that the optimal pairing momentum $\bm{Q}(\eta)$ is independent of the temperature, resulting in a horizontal boundary line between the BCS and FFLO phases. In this case, the thermal fluctuations play a negligible role and the phase transition is solely driven by QGD. The calculated $T_{c0}(\eta)$, $T_{c\bm{Q}}(\eta)$ and $\bm{Q}(\eta)$ exhibit quantitative agreement with the corresponding analytical results in Fig. 3(a).
@@
To elucidate the nature of the phase transition, in Fig. 4(c) to (e), we show the landscapes of free energy $F(\bm{q})$ at the three points (blue, green, and pink dots) in Fig. 4(a). Initially, in the BCS region ($\eta<\eta_{c}$) [Fig. 4(c)], the global minimum of the free energy $F(\bm{q})$ is at $\bm{q}=0$, which is stable under weak perturbations. Some metastable states exist at $\bm{q}\neq 0$ as local minima in $F(\bm{q})$. When $\eta>\eta_{c}$ [Fig. 4(d)], $\bm{q}=0$ turns into a local minimum but remains a metastable state, while some $\bm{q}=\bm{Q}$ points become global minima (the model respects $C_{4}$ rotational symmetry, so the minima have four-fold degeneracy in the whole first Brillouin zone, see SI III.A sup ), the FFLO phase can manifest. This indicates that the QGD energetically penalizes BCS pairing more than FFLO pairing. A first-order phase transition from BCS to FFLO occurs near $\eta_{c}$, which is consistent with the results obtained from the pairing susceptibility. As $\eta$ continues to increase [Fig. 4(e)], the local minimum at $\bm{q}=0$ eventually nearly disappears.
VII Robustness with finite bandwidth
The QGD-driven BCS-FFLO phase transition has been examined in the perfect flat-band limit, both analytically and numerically. To show the robustness of QGD-induced BCS-FFLO transition with finite band dispersion, we modify the Bloch Hamiltonian in Eq. (S30) as $\mathcal{H}^{\prime}_{\bm{k}\sigma}=\mathcal{H}_{\bm{k}\sigma}-W\left(\cos k_{x}+\cos k_{y}\right)\lambda_{0}/4$, where an additional nearest hopping term brings a finite bandwidth $W$ as shown in Fig. 2(b). In this case, the Bloch wave functions remain unchanged. We keep $W\ll U$ to ensure a narrow bandwidth compared to the interaction.
@@ VII Robustness with finite bandwidth
In Fig. 4(b), the mean-field phase diagram at $W/U=0.12$ ($U=0.4t$) is shown. Self-evidently, compared with Fig. 4(a) ($W/U=0$), the transition boundary line is temperature-dependent and becomes curved due to the nonzero $W$, requiring a larger QGD to reach the FFLO region, especially at lower temperatures. This fact indicates that increasing the temperature may also drive the BCS-FFLO phase transition [e.g., along the horizontal dashed line at $\eta=0.25$ in Fig. 4(b)]. The reason is, in Eq. (5), the form factor gets Fermi-surface averaged for a dispersive band, which weakens the effects of QGD. And the thermal excitation energy $k_{B}T$ influences the Fermi-surface average in Eq. (5), reshaping the boundary line. On the other hand, for conventional superconductors with a highly dispersive band, we have $W\gg U\gg k_{B}T$ in the weak coupling regime, where the QGD-driven mechanism for the FFLO state may not be effective (more discussions, see SI II sup ). Interestingly, previous studies of spin-imbalanced Fermi gases found that the FFLO state becomes unfavorable when the pairing interaction is sufficiently strong Radzihovsky and Sheehy (2010), which stands opposite to the QGD case.
Figure 5: (a) 1D SSH-like model with nearest-neighbor hopping $u$ and $v$, second nearest-neighbor hopping $t_{a}$ and $t_{b}$, and third nearest-neighbor hopping $w$. (b) The band structure of the model when $\phi=0$ (black solid line) and $\phi=0.1\pi$ (red dashed line). (c) The distributions of quantum metric $g(k)$ at $\phi=0$, $0.01\pi$, and $0.05\pi$ (labeled in black, red, and blue lines, respectively). (d) The critical temperatures of the most favored FF states and the BCS state are plotted as functions of $\phi$ in red and blue lines, respectively. The results from self-consistent mean-field (MF) calculations are denoted by the solid lines, and the analytical results from Eq. 6 (PS) are denoted by the dashed lines. (e) The favored finite-momentum $Q$ with varying $\phi$. Parameters used here: $v=|w|=t_{a}=t_{b}=0.4t,u=-t,U=0.1t,\nu=0.5.$
VIII Flat-band FF state with inversion breaking
While we did not deliberately distinguish the FF and LO pairings throughout previous sections, we expect the LO-type pairing state to be more stable than the FF type due to the $\pm\bm{Q}$ degeneracy (see Fig. 4) for the inversion-symmetric model [Eq. (S30)]. Nevertheless, in the following we further demonstrate that if the inversion symmetry is broken in the quantum metric rather than band dispersion, the $\pm\bm{Q}$ degeneracy will be lifted, resulting in the FF-type pairing state driven by QGD.
Here we consider a 1D spinless SSH-like model, as shown in Fig. 5(a), with the complex nearest-neighbor hopping terms $u$ and $v$, and third nearest-neighbor hopping $w$ Arora et al. (2022). We add the next nearest-neighbor hopping terms $t_{a}$ and $t_{b}$ to flatten the lower band [see Fig. 5(b)], then Hamiltonian can be written as
$H_{1\mathrm{D}}(k)=\left(\begin{array}[]{cc}2t_{a}\cos(k)-\mu&J^{*}(k)\\ J(k)&2t_{b}\cos(k)-\mu\end{array}\right),$ (11)
where $J(k)=|u|e^{-i(k+\phi/2)/2}+|v|e^{i(k+\phi/2)/2}+|w|e^{-i(3(k+\phi/2)/2-\phi)}$, and the additional phase shift $\phi/2$ of $k$ is added to make the band dispersion symmetric, i.e., the group velocities satisfy $v_{G}(\bm{k})=-v_{G}(-\bm{k})$.
In Fig. 5(c), we can observe that the quantum metric is asymmetric as $g(k)\neq g(-k)$ when $\phi\neq 0$, which violates the inversion symmetry. At the same time, QGD renders $g_{\uparrow}(k)\neq g_{\downarrow}(-k)$ when we introduce the spin degree of freedom. Therefore, the superconducting BdG Hamiltonian can be expressed as
$H_{\mathrm{BdG}}(k,q)=\left(\begin{array}[]{cc}H_{1D}(k+q)&\hat{\Delta}_{q}\\ \hat{\Delta}_{q}^{\dagger}&-H_{1D}^{T}(-k)\end{array}\right),$ (12)
where $\hat{\Delta}_{q}=\mathrm{diag}[\Delta_{A}(q),\Delta_{B}(q)]$. Following the similar procedures as the previous sections, we can find the critical temperatures as well as the finite-momentum $Q$ of the most favored FF pairing state at the flat band filling, from the self-consistent mean-field (MF) calculations [Fig. 5(d)], and from the pairing susceptibility (PS) analysis Eq. 6 [Fig. 5(e)]. Interestingly, in Fig. 5(d), the critical temperature of the most favored FF state will immediately become higher than the BCS state when $\phi$ is turned on. And in Fig. 5(e), the finite-momentum $Q$ of the most favored FF state is an odd function of $\phi$. Besides, we can see that the mean-field results basically match well with the predictions from Eq. 6, which applies in principle only at the exact flat-band case. Due to the fact that the lower band of this model is not exactly flat when $\phi\neq 0$, and becomes more and more dispersive with increasing $\phi$ [see Fig. 5(b)], the numerical and analytical results slightly deviate from each other when $\phi$ increases.
The characteristics in Fig. 5(d) and (e) are quite different from the QGD-driven FFLO phase for the $\zeta$-lattice model in Fig. 3 and Fig. 4, where the system experiences a first-order BCS-FFLO phase transition driven by QGD. Nevertheless, they mimic the feathers of helical superconductors with spin-orbit interaction Smidman et al. (2017); Kinnunen et al. (2018). In SI III.B sup , we provide another scheme to realize the flat-band FF pairing state and the emergent superconducting diode effect by turning on an extra inversion-breaking term in Eq. (S30).
IX Conclusion and Discussion